write down the expression of critical pressure ,critical volume and critical temperature in terms of wander waal`constants?

Grade:12th pass

1 Answers

Umakant biswal
7 years ago
@ shravan 
as per the van der walls equation 
crtical temp =  8a / 27bR 
critical pressure = a/ 27 b square 
where a and b are constants 
oir simply u can say 
P = RT / V/(n-b) = a/ (v/n) square 
hope it clears 
all the best